In G.P the first term is 153 and the sixth term is 17/27 find the sum of the first four terms
Answer by Frosty7(4)

T1 = 153
T6 = 17/27
Equation for nth term is ar^(n-1)
So, T1: 153r^0=153 ----- (1)
T6: 153r^5= 17/27 ------ (2)
(2) ÷ (1)
Which gives us, r^5=1/243
r = 1/3
Sum of n terms is Sn= a((1-r^n)/(1-r))
So, S4= 153((1-〖1/3〗^4)/(1-1/3))
=680/3
Hope this helps!